Solve the system by substitution.
x – 2y = 19
2x + 5y = –7

Answer by LilSkittleMd(119)

x – 2y = 19
2x + 5y = –7
First solve one equation for a variable
x = 2y + 19
Now plug it into the x value in the second equation
2(2y + 19) + 5y = -7
Use the distributive property
4y + 38 + 5y = -7
9y + 38 = -7
Subtract 38 from both sides of the equation
9y = -45
Divide both sides of the equation by 9
y = -5
Now plug the y value into one of the original equations
x - 2(-5) = 19
x + 10 = 19
x = 9